OpenStudy (anonymous):

int(x^2*sqrt(x-1), x = 1 .. 2)

hartnn (hartnn):

do the substitution u=x-1

OpenStudy (anonymous):

making du=dx

OpenStudy (anonymous):

so you get int(x^2*sqrt(u) ...which to me is no help

hartnn (hartnn):

as u=x-1, x = u+1 you will change x^2 to (u+1)^2 too :) now try..

OpenStudy (anonymous):

oh lol.....thank you! just having that moment!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

well..wait...ha....i am getting (u+1)^2*sqrt(u) ..i can't simplify anymore to get a known integral

OpenStudy (loser66):